OpenStudy (anonymous):

A magic show sold tickets for $2 for every child and $3 for every adult. The total sales were $118. The difference of two times the number of adults and half the children will total 20. Using the elimination method for the system of equations, what is the number of adults and children at the magic show? A. adults = 5 children = 30 B. adults = 18 children = 32 C. adults = 16 children = 35 D. adults = 32 children = 18

OpenStudy (anonymous):

3a + 2c = 118 -->(2) 2a - .5c = 20 -->(-3) ------------ 6a + 4c = 236 -6a + 1.5c = - 60 --------------add 5.5c = 176 c = 176/5.5 c = 32 3a + 2c = 118 3a + 2(32) = 118 3a + 64 = 118 3a = 118 - 64 3a = 54 a = 54/3 a = 18 adults 18, children 32

OpenStudy (anonymous):

You did a great job :) at explain

OpenStudy (anonymous):

sorry I took so long....I kept messing up ...lol. I finally got it :)

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

you did great

OpenStudy (anonymous):

thanks :)